find the number of distinct positive integers that can be formed using 0,1,2,3 where each integer is used at the most once.

Grade:11

1 Answers

arpit burad
13 years ago

1 digit integer=4

2 digit integer=4p2-3p1=9

3 digit integer=4p3-3p2=18

4 digit integer=4p4-3p3=18

adding these we get 49 which is the answer